OpenStudy (anonymous):

Group Theory Question. Can a right coset strictly contain a left coset? More precisely, is there a group \(G\), with subgroup \(H\), and an element \(g \in G\), such that \(gH \subseteq Hg\) but \(gH \neq Hg\)? I understand the answer is clearly "no" if G is finite, since |gH| = |Hg| always, so then \(gH \subseteq Hg\) clearly implies \(gH = Hg\). But for infinite groups this is not enough... Thanks!

OpenStudy (anonymous):

@zzr0ck3r @KingGeorge The only people on here who I know post group theory questions... :)

OpenStudy (anonymous):

would have to come with a non abelian infinite group where \(gH\subset Hg\)

OpenStudy (anonymous):

but i can't think of one off the top of my head where containment is strict

OpenStudy (anonymous):

matrices come to mind

OpenStudy (anonymous):

Indeed. And stay away from normal subgroups altogether... I thought perhaps \(GL_n(\mathbb{R})\) might be a good place to start looking. Not sure how to start looking though! I was rather hoping someone might know some facts which would prove that the answer was no.

OpenStudy (anonymous):

i think we can come up with an example, may take a minute ( or more )

OpenStudy (anonymous):

\[H=\left[ {\begin{array}{cc} a & 1 \\ 0 & 1 \\ \end{array} } \right]\] maybe? and \(g=\left[ {\begin{array}{cc} 2 & 0 \\ 0 & 1 \\ \end{array} } \right]\)

OpenStudy (anonymous):

maybe not but i actually remember something like this with linear maps \(G\) is the set of linear maps \(g(x)=ax+b\)

OpenStudy (anonymous):

yeah the gimmick is to make \(H\) be the subgroup where \(b\) is an integer hold on a second

OpenStudy (anonymous):

Linear maps? Is the group operation composition?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

i forget exactly how it works though

OpenStudy (anonymous):

some how you get something is an integer and something isn't so containment is strict

OpenStudy (anonymous):

i know \(H\) is the set where \(h(x)=ax+n, n\in m\mathbb{Z}\)

OpenStudy (anonymous):

the m was a typo, \(H\) is the set of elements \(h(x)=ax+n,n\in \mathbb{Z}\)

OpenStudy (anonymous):

That makes sense actually. I was hoping to get something with integers and something with only even integers or something.

OpenStudy (anonymous):

ok lets take \(g\) as \(g(x)=2x+3\) maybe that will work

OpenStudy (anonymous):

lol thank god for google here is a much better explanation see example 2 here http://math.stackexchange.com/questions/460247/example-of-a-group-with-non-equal-coset-for-a-fixed-element

OpenStudy (anonymous):

Wow. You put my googling abilities to shame...

OpenStudy (anonymous):

wish they had had this when i was in school would have 2 or 3 PhD by now

OpenStudy (anonymous):

Ah, but where's the fun when you can just google the answer instead of working things out? ;-)

OpenStudy (anonymous):

well we have the satisfaction at least of knowing we were on the right track examples like this are usually canonical examples, the one or two you know but probably would not find on your own without significant hints

OpenStudy (anonymous):

how many continuous nowhere differentiable functions do you know? or uncountable sets of measure zero?

OpenStudy (anonymous):

I suppose so. I wonder how people came up with these. Experimentation? Experience? I got as far as looking for a non-abelian infinite group. Everything after that seemed to be "try some well-known groups and hope for the best"!

OpenStudy (anonymous):

my guess is either none or one

OpenStudy (anonymous):

you get the book called "counter examples in analysis" or "counter examples in algebra"

OpenStudy (anonymous):

I've heard of the Wierstrass function (or however its spelt). That's all I know. As for the second question, I don't know what "measure zero" means!

OpenStudy (anonymous):

find the thing that "is this but not that", proving that the two sets are not equivalent

OpenStudy (anonymous):

you will know it soon enough i am sure means has no length basically or its probability is zero the example you will know will be called the cantor set, and the only other ones you will know will be variations on that theme have fun

OpenStudy (anonymous):

Sounds interesting! I look forward to (second year) maths next term... Is there a book "Counterexamples in Algebra"? I've seen a "Counterexamples in Analysis" book, but not for algebra.